### v2

@@ 8 Conclusions and Further Work
In this exploratory paper, we have presented a BPMN extension to model human-agentic workflows. Our extension enables the modeling of complex collaboration patterns between humans and agents, including specifying the reflection strategies agents can do on specific tasks.
Nevertheless, to provide a more complete framework for human-agentic systems, we want to explore the following lines of future work. At the modeling level, we plan to provide a sublanguage to specify in detail the governance and decision-making strategies as part of the merging nodes for agents’ collaboration efforts. This language would allow us to go from general strategies (e.g. voting) to a much more refined and unambiguous specification (e.g. how the voting needs to be precisely evaluated and what conditions must be met to decide the result and implications of the voting).
We also want to propose an uncertainty propagation mechanism that, given the overall uncertainty of the agents and their confidence in the result of a given task, assigns an overall uncertainty of the task, which should then be propagated also to the consecutive tasks. This value could be used in the gateways to ensure a certain level of trustworthiness in the decision-making process (e.g. automatically going back to previous tasks when the confidence level is not over the threshold).
@@
Additionally, we plan to add a code generator to the tool support able to produce an executable representation of the model, including the governance aspects and the uncertainty propagation mechanisms. The runtime engine could be, potentially, a combination of BPEL engines and agentic platforms, where the BPEL engine would take care of the global orchestration and delegate to the agentic platforms the execution of the agent tasks.
Finally, we plan to empirically evaluate our extension and explore its application in industrial use cases.
{credits}
8.0.1 Acknowledgements
This work is part of the project TED2021-130331B-I00 funded by MCIN/AEI/10.13039/501100011033 and European Union NextGenerationEU/PRTR; and BESSER, funded by the Luxembourg National Research Fund (FNR) PEARL program, grant agreement 16544475.

### v3

@@ VIII Conclusions and Further Work
We have presented a BPMN extension to model human-agentic workflows. Our extension enables the modeling of complex collaboration patterns between humans and agents, including specifying the agents’ reflection strategies.
As further work, we plan to provide a sublanguage to specify in detail the governance and decision-making strategies as part of the merging nodes for agents’ collaboration efforts (e.g., should decisions be based on consensus? voting strategies…). We will also propose an uncertainty propagation mechanism that, given the overall uncertainty of the agents and their confidence in the result of a given task, assigns an overall uncertainty of the task, which should then be propagated also to the consecutive tasks.
@@
Additionally, we plan to work on code generators aimed at producing an executable representation of the model, including the governance aspects and the uncertainty propagation mechanisms to make operational the modeled workflows. The generators could target, potentially, a combination of BPEL engines and agentic platforms, where the BPEL engine would take care of the global orchestration and delegate to the agentic platform the task execution.
Finally, we plan to empirically evaluate our extension and explore its application in industrial use cases. In particular, to ensure the cognitive effectiveness of our notation, we plan to conduct a metric-based assessment based on the alignment with Moody’s principles for visual notations [10], following our previous work on collaborative modeling notation evaluation  [1].